Hi
At a seminar 2/5 of the audience were students. 3/10 of them were retirees the rest were professionals. There were 18 more students than professionals.
How many were at the seminar.
Thanks

Answer by ikleyn(52776)

.
Let x be the number of students (the entire audience).


Then from the condition


    0.4x - 0.3x = 18.


or


       0.1x    = 18.


Hence,


          x    = 18/0.1 = 180.


ANSWER.  The question is  "How many were at the seminar", but it does not specify "how many of whom".


         The entire audience was 180 people, in all.